# Inhomogeneous mixing: From microscopic dynamics to mesoscopic staircases

### v1

@@ 3.2 Observations of the $E\times B$ staircase
In KSTAR L-mode plasmas [42, 79], the two dimensional radial corrugation of electron temperature was measured, appearing as concentric jet-like patterns in the normalized temperature images in Figure 8(b). It was only observed in the phase where the MHD instabilities were suppressed and ballistic heat flux propagation events of various sizes, similar to avalanches in the near-marginal regime, are dominant [42, 79]. With help of a large diagnostic view, the dynamics of the temperature corrugation could be investigated in KSTAR plasmas. These transport barriers are intermittent but persistently observed. Their ‘typical’ lifetime is on the order of milliseconds, but the distribution of lifetime has a long tail [79]). They can drift radially and their tread width can change in time [42, 79]. It was also found that they can impede the ballistic propagation of the heat flux until they are dissipated [42, 79], demonstrating that this structure indeed corresponds to transport barriers. The radial correlation length of density fluctuations was also limited within the mixing zone width ($\Delta$[79].
Figure 9: Transfer entropy $\mathrm{TE}(\rho,\tau)$ measured in (a) the TJ-II experiment and (b) the W7-X experiment. Reprinted from Milligen et al. [80] (Copyright 2017 IOP) and Milligen et al. [81] (Copyright 2018 IOP), respectively.
@@
The third approach is to trace the propagation of spontaneous pulses that are sufficiently mild not to perturb transport barriers. In TJ-II [80] and W7-X [81] L-mode plasmas, the propagation of spontaneous and mild electron temperature perturbations was analyzed using an information theoretic measure called transfer entropy. The perturbations were too irregular (as avalanches) for their propagation to be analyzed using the conventional Fourier technique. Transfer entropy $\mathrm{TE}(\rho,\tau)$ measures the causal relation or the information transfer between two serial data separated by $\rho$ in space and $\tau$ in time, based on the degree of enhanced predictability by knowing the history of additional data [82]. The analysis identified the trapping (barrier) and jumping (mixing) zones of the information contained in a reference signal measured near the plasma center (see Figure 9). Most barriers were found around the low-order rational surfaces, and they were interpreted as a result of the poloidal flow shear associated with the rational surfaces [83]. The immediate transfer of the information in the jumping zone (indicated by red arrows) was attributed to the interacting modes between different rational surfaces. The observed inhomogeneous mixing of the information in TJ-II and W7-X plasmas suggests the importance of electromagnetic effects or non-adiabatic electron effects on a plasma state with mesoscopic transport barriers that could not be investigated in the electrostatic and adiabatic electron gyrokinetic simulations. Barriers from different mechanisms may be not necessarily exclusive.
3.3 Characteristics of the $E\times B$ staircase
@@ 3.3.1 The lifetime
The $E\times B$ staircase can undergo merging, splitting, drifting, dissipation, and even destruction and reformation with avalanches [84, 67, 42, 79]. In KSTAR experiments, the lifetime $t_{\mathrm{life}}$ of the $E\times B$ staircase was estimated by analyzing the high resolution spatio-temporal dynamics of the two-dimensional temperature corrugation [^1] In other experiments the measurement of the lifetime was technically impossible. The fast sweeping reflectometry used in Tore Supra plasmas [69, 76] was limited to capture the instantaneous moments of the $E\times B$ staircase. In HL-2A, DIII-D, TJ-II, and W7-X experiments, about 3 milliseconds (HL-2A [77]), 40 milliseconds (DIII-D [78]), or 50—200 milliseconds (TJ-II [80] and W7-X [81]) records of the diagnostics data were utilized to identify the (on average) transport barriers during the analysis periods.. Here, the lifetime was defined as the duration from the initial appearance through all its deformations (merging or splitting) until its disappearance. It appears to dissipate alone or be penetrated by large avalanches (see Figure 4 of reference [79]), having the lifetime of the order of 0.1—1.0 milliseconds [79]. Although the range of measured lifetimes was too narrow to conclude, it seems to follow a long tail distribution which can be fit with a two-fold power-law function [79]. Note that distributions of characteristics of self-organized structures in non-equilibrium complex systems are rarely normal.
The lower ion collisionality seems to be favorable to the robustness of transport barriers. A weak and scattered trend of decreasing lifetime with respect to the collisionality increase was observed within the banana regime $\nu_{*}<1$ in KSTAR [79]. In Tore Supra experiments, the $E\times B$ staircase was observed mostly in the banana regime [67]. Note that in the GYSELA simulations the staircase is only weakly affected by the collisionality in the banana regime and it starts to be damped around the plateau transition [67].
@@ 3.3.2 The barrier strength
The barrier strength was defined as the integrated zonal mean shear strength over the flow layer thickness $\delta^{\mathrm{flow}}$ in the GYSELA simulation [67]. It indicates how impermeable the flow shear layer is to the radial transport or mixing processes. Since the flow shear is hard to be measured accurately in experiments, a proxy of the barrier strength was measured. In Tore Supra experiments [76], the height and the width of the correlation length reduction were measured. Since both quantities decrease with $\rho_{*}^{-1}$, the barrier would be weaker as the plasma size increases. In HL-2A experiments [77], the permeability was estimated by how the zero-time delay cross correlation between radially adjacent channels varies with the normalized fluctuation amplitude. The regions of enhanced and suppressed permeability (mixing) were identified. The enhanced permeability region coincides with the larger correlation length region, and the suppressed permeability regions, characterized by the relatively low correlation coefficients for the smaller amplitude fluctuations, are located near the the correlation length minima or the pressure gradient maxima as expected. In KSTAR experiments [79], the square of temperature corrugation amplitude was measured. They exhibit a skewed distribution, and the larger amplitude corrugation was less observed. The squared amplitude seems to have a weak and scattered decreasing trend with respect to the collisionality increase.
@@ 3.3.3 The tread width
The tread width $\Delta$ of the $E\times B$ staircase, or the mixing zone width, corresponds to the (maximum possible) radial scale of avalanche transport. In Tore Supra experiments [67], $\Delta$ was taken to be the distance between the minima of the correlation length profile, while in KSTAR experiments [79], it was defined as the inverse of the average radial wavelength of the temperature corrugation. The scaling and the distribution of $\Delta$ was investigated in Tore Supra and KSTAR experiments, respectively, based on the large number of measurements.
@@
The scaling of $\Delta$ with respect to $\rho_{*}=\rho_{s}/a$ has been of particular interest since it can be related to the global confinement scaling. In the GYSELA simulations [67], $\Delta$ does not seem to simply increase as $\rho_{*}^{-1}=a/\rho_{s}$ increases as shown in Figure 10(a). Instead, the most probable value has been fixed around $40~\rho_{s}$ for investigated values of $\rho_{*}$. This feature expects the favorable gyro-Bohm scaling. However, in Tore Supra experiments, $\Delta$ seems to increase with $\rho_{*}^{-1}$ as shown in Figure 10(b), implying that even with the avalanche regulation by the $E\times B$ staircase, the global confinement scaling can be like the Bohm scaling [76]. Further research would be required to address this inconsistency.
Figure 10: (a) The tread width $\Delta$ versus $\rho_{*}^{-1}$ in the GYSELA simulation. Reprinted from Dif-Pradalier et al. [67]. Copyright 2017 IOP. (b) $\Delta$ versus $\rho_{*}$ and small and large $\Delta$ measurements in the Tore Supra experiment. Reprinted from Hornung et al. [76]. Copyright 2017 IOP. The distribution of $\Delta$ in (c) the GYSELA simulation and (d) the KSTAR experiment. Reprinted from Dif-Pradalier et al. [67] (Copyright 2017 IOP) and Choi et al. [79] (Copyright 2024 IOP), respectively.
@@
The distribution of $\Delta$ ($P(\Delta)\sim P_{\text{Fréchet}}$) was found to follow a long tail distribution called the Fréchet distribution in the GYSELA simulations [67] and also in KSTAR experiments [79] as shown in Figures 10(c) and (d). The Fréchet distribution is one of the extreme value distributions, a distribution of maxima of samples when the samples follow a long tail distribution such as the power-law distribution. This Fréchet distribution can imply dual roles of avalanches, co-existing ballistic transport events (mixing process) as explained in reference [79]. They can destruct the $E\times B$ staircase but may also contribute to its formation and $\Delta$ determination [^2] The distribution of temperature perturbation size of avalanches ($S\sim\delta T^{2}$ [79]) was found to follow a power-law distribution ($P(S)\sim P_{\mathrm{power-law}}$[42]. Assuming that (1) a temperature perturbation beyond some threshold can trigger the formation of the $E\times B$ staircase and (2) $\Delta$ depends on the size of the triggering temperature perturbation which is rare and whose size would be the maximum size of perturbations for some period, the Fréchet distribution of $\Delta$ follows from its definition ($P(\Delta\sim\max\{S\})\sim P_{\text{Fréchet}}$). Some models have shown these behaviors, e.g., the threshold [85] and the turbulence spreading strength $\sim\Delta$ [86]. Also, note that there are minor temperature corrugations that form away from large avalanches without apparent interactions, and their $\Delta$ distribution does not seem to follow a Fréchet distribution [79]. .
@@ 4 Summary and discussion
    There exist strong interactions between blobs/voids and $E\times B$ sheared flow. The turbulence spreading induced by blobs decreases as the flow shearing rate increases, while the turbulence spreading induced by voids is insensitive to the shearing rate. On the other hand, voids can excite drift waves and so drive intermittent zonal flow.
@@
The $E\times B$ staircase, or more generally mesoscopic transport barriers, represents a global realization of inhomogeneous mixing in fusion plasmas. Various methods have been utilized to identify these elusive transport barriers in fusion core plasmas. As summarized in Table 3, they have a lifetime order of 0.1–1.0 ms with a long tailed distribution (KSTAR) and a tread width of 10–100 $\rho_{s}$ (Tore Supra, HL-2A, TJ-II, W7-X) with a Fréchet distribution (KSTAR). Their strength seems to decrease with $\rho_{*}^{-1}$ (Tore Supra) or $\nu_{\mathrm{i}}$ (KSTAR), and they seem to be associated with the low-order rational surfaces (HL-2A, KSTAR, TJ-II, W7-X).
Although some progress has been made over the past decade in the experimental side, a comprehensive understanding of the physics of the $E\times B$ staircase remains elusive. Various models have provided different explanations for the $E\times B$ staircase (see Appendix), which reflects how profound the problem is. Most models may be classified as variants of a few basic types, listed below.
@@
    Bistable mixing – in which a system manifests multiple mixing lengths – is a natural route to layering. Frequently, one such length is emergent and intensity dependent – such as the Rhines, Ozmidov or Hinze scales. In such cases, the flux-gradient relation becomes nonlinear. In magnetized plasma, zonal flow generation can produce bistability, and lead to transport barrier formation. Ultimately, mesoscopic bistability has its origins in the microphysics of the inhomogeneous mixing.
@@
    Homogenization, in which localized cellular mixing sharpens inter-cellular PV gradients (i.e. shear sharpening), is a natural way for an array of cells to induce a layered profile structure. Homogenization requires some separation of time scales, i.e. $\tau_{turn}<\tau_{mix}<\tau_{diff}$, so a cell turn over time is shorter than a mixing time which, in turn, is shorter than a cell diffusion time. Homogenization is, in some sense, the simplest mechanism, and requires only and array of cells with clear bounding separatrices.
@@
    Phase separation – as in spinodal decomposition – is a natural layering mechanism. By their nature negative viscosity phenomena may be classified as phase separation processes. Jamming can trigger phase separation. In MFE (magnetic confinement energy), heat flux jams can trigger temperature staircase formation.
The results of recent blobs/voids research can provide helpful insights into developing a better physical model for the $E\times B$ staircase (avalanches and zonal flow layers). Incorporating turbulence intensity flux and its dynamics into physical models of the $E\times B$ staircase is necessary. In fact, such models have been developed but have not yet been exercised effectively. In particular, the length scale for the turbulence intensity can differ from that for the transport flux, adding a scale to the bistable mixing model [86]. This is plausible, since research on coherent structures [7] indicates a departure of the spreading flux from the particle flux. The scale of the spreading flux effectively sets the scale of the turbulence envelope, which in turn determines the zonal flow. Also, Figure 4 shows that the symmetry breaking of turbulence intensity flux, possibly related to the turbulence intensity gradient [73], yields a net turbulence spreading and forms a region of enhanced mixing. Note that, in previous avalanche models [70] or simulations [74], the variation of the pressure gradient or the associated $E\times B$ shear by blobs/voids was responsible for a directional propagation of blobs/voids. On the other hand, the interaction between blobs/voids and zonal flows implies that the turbulence influx through the separatrix, which can drive zonal flows in the core plasma [87, 88, 9], may contribute to the formation or sustainment of the $E\times B$ staircase. The correlated growth of zonal flows at distant regions might be mediated by turbulence spreading between zonal flow layers [89].
A key question for the $E\times B$ staircase is the scale selection problem – i.e. what sets the separation between zonal flow layers? There are several possibilities:
@@
    Bistable mixing (2 scales model). Here, the emergent scale adjusts to define the layer separation. This is an outcome of the nonlinear flux-gradient relation. In the case of drift-Rossby turbulence, the Rhines scale is selected.
@@
Further research – experimental, theoretical and computational – is necessary to sort out which mechanism is relevant. Experimental research includes, but not limited to, (1) the multi-machine analysis, (2) the development of a stable staircase scenario (possibly with the help of energetic particles [90, 91] or light impurities [92]), and (3) the development of fine-scale flow or potential diagnostics, which enables the detailed analysis of inhomogeneous mixing process (see Section 2), in the core plasma.
@@
4.1 Models on the $E\times B$ staircase

### v2

@@ 3.2 Observations of the $\mathbf{E}\times\mathbf{B}$ staircase
In KSTAR L-mode plasmas [42, 80], the two-dimensional radial corrugation of electron temperature was measured as shown in Figure 8(b). It was observed in the phase where the MHD instabilities were suppressed and ballistic heat flux propagation events of various sizes, similar to avalanches in the near-marginal regime, are dominant [42, 80]. The dynamics of the temperature corrugation could be investigated using the high resolution image data, and the statistics of its characteristics could be analyzed [80] (see below). These transport barriers are intermittent but observed persistently. Barriers can drift radially and their tread width can change in time [42, 80]. Also, they can impede the ballistic propagation of the heat flux until they are dissipated [80], demonstrating their capability as transport barriers.
Figure 9: Transfer entropy $\mathrm{TE}(\rho,\tau)$ measured in (a) the TJ-II experiment and (b) the W7-X experiment. Horizontal thin dashed lines indicate the locations of the rational surfaces of the corresponding $q$ values. Thick dashed lines indicate the propagation/trapping of the information revealed by $\mathrm{TE}(\rho,\tau)$. Red arrows indicate the immediate transfer of the information. Reprinted from Milligen et al. [81] (Copyright 2017 IOP) and Milligen et al. [82] (Copyright 2018 IOP), respectively.
@@
The third approach is to trace the propagation of spontaneous pulses that are sufficiently mild so not to perturb transport barriers. In TJ-II [81] and W7-X [82] L-mode plasmas, the propagation of spontaneous and mild electron temperature perturbations was analyzed using an information theoretic measure called transfer entropy. The perturbations were too irregular for their propagation to be analyzed using the conventional Fourier technique. Transfer entropy $\mathrm{TE}(\rho,\tau)$ measures the causal relation or the information transfer between two serial data separated by $\rho$ in space and $\tau$ in time, based on the degree of enhanced predictability by knowing the history of additional data [83]. The analysis identified the trapping (barrier) and jumping (mixing) zones of the information (see Figure 9). Most barriers were found around the low-order rational surfaces, and they were interpreted as a result of the poloidal flow shear associated with the rational surfaces [84]. The immediate transfer of the information in the jumping zone (indicated by red arrows) was attributed to the interacting modes between different rational surfaces. A link between the rational surfaces and transport barriers suggests the importance of electromagnetic effects or non-adiabatic electron effects (see below). Barriers from different mechanisms may be not necessarily exclusive.
3.3 Characteristics of the $\mathbf{E}\times\mathbf{B}$ staircase
@@ 3.3.1 The lifetime
The $\mathbf{E}\times\mathbf{B}$ staircase can undergo merging, splitting, drifting, dissipation, and even destruction and reformation with avalanches [85, 68, 42, 80]. In KSTAR experiments, the lifetime $t_{\mathrm{life}}$ of the $\mathbf{E}\times\mathbf{B}$ staircase was estimated by analyzing the high resolution spatio-temporal dynamics of the two-dimensional temperature corrugation [^2] In other experiments, the measurement of the lifetime was technically impossible. The fast sweeping reflectometry used in Tore Supra plasmas [70, 77] was limited to capture the instantaneous moments of the $\mathbf{E}\times\mathbf{B}$ staircase. In HL-2A, DIII-D, TJ-II, and W7-X experiments, about 3 milliseconds (HL-2A [78]), 40 milliseconds (DIII-D [79]), or 50—200 milliseconds (TJ-II [81] and W7-X [82]) records of the diagnostics data were utilized to identify the (on average) transport barriers during the analysis periods.. Here, the lifetime was defined as the duration from the initial appearance through all its deformations (merging or splitting) until its disappearance. It appears to dissipate or be penetrated by large avalanches (see Figure 4 of reference [80]), having the lifetime of the order of 0.1—1.0 milliseconds [80]. Although the range of measured lifetimes was too narrow to conclude, the staircase lifetime seems to follow a long tail distribution which can be fit with a two-fold power-law function [80]. Note that distributions of characteristics of self-organized structures in non-equilibrium complex systems are rarely normal.
The lower ion collisionality seems to be favorable to the robustness of transport barriers. A weak and scattered trend of increasing lifetime with respect to the collisionality decrease was observed within the banana regime $\nu_{*}<1$ in KSTAR [80]. In Tore Supra experiments, the $E\times B$ staircase was observed mostly in the banana regime [68]. Note that in the GYSELA simulations the staircase is only weakly affected by the collisionality in the banana regime and it starts to be damped around the plateau transition [68].
@@ 3.3.2 The barrier strength
The barrier strength was defined as the integrated zonal mean shear strength over the flow layer thickness $\delta^{\mathrm{flow}}$ in the GYSELA simulation [68]. It indicates how impermeable the flow shear layer is to the radial transport or mixing processes. Since the flow shear is hard to be measured accurately in experiments, a proxy of the barrier strength was measured. In Tore Supra experiments [77], the height and the width of the correlation length reduction were measured. Since both quantities decrease with $\rho_{*}^{-1}=a/\rho_{s}$, the barrier would be weaker as the plasma size increases. In HL-2A experiments [78], the permeability was estimated by how the zero-time delay cross correlation between radially adjacent channels varies with the normalized fluctuation amplitude. The regions of enhanced and suppressed permeability (mixing) were identified. The enhanced permeability region coincides with the larger correlation length region. The suppressed permeability regions are located near the correlation length minima or the pressure gradient maxima as expected. In KSTAR experiments [80], the square of temperature corrugation amplitude was measured. This exhibited a skewed distribution, and the larger amplitude corrugation was less frequently observed. The squared amplitude seems to have a weak and scattered increasing trend with respect to the collisionality decrease.
@@ 3.3.3 The tread width
The tread width $\Delta$ of the $E\times B$ staircase, or the mixing zone width, corresponds to the (maximum possible) radial scale of avalanche transport. In Tore Supra experiments [68], $\Delta$ was taken to be the distance between the minima of the correlation length profile. In KSTAR experiments [80], it was defined as the inverse of the average radial wavelength of the temperature corrugation. The scaling and the distribution of $\Delta$ were investigated in Tore Supra and KSTAR experiments, respectively, based on the large number of measurements.
Figure 10: Various measurements of the tread with $\Delta$. It is called step size in (a) and (c), avalanche (spatial) scale in (b) and staircase (tread) width in (d). (a) The normalized tread width $\Delta/\rho_{s}$ versus $\rho_{*}^{-1}=a/\rho_{s}$ where $\rho_{s}$ is the ion sound gyroradius in the GYSELA simulation. Reprinted from Dif-Pradalier et al. [68]. Copyright 2017 IOP. (b) $\Delta/\rho_{s}$ versus $\rho_{*}$, and the radial correlation length $\langle L_{coh}\rangle_{F}$ profile measurements for small and large $\Delta$ cases in the Tore Supra experiment. $r_{s_{1}}$ is the radial location of the first (expected) shear flow layer or the first $\langle L_{coh}\rangle_{F}$ minimum. Reprinted from Hornung et al. [77]. Copyright 2017 IOP. The distribution of $\Delta$ in (c) the GYSELA simulation and (d) the KSTAR experiment. Here, $\kappa$, $\Delta_{\mu}$ and $\sigma$ are shaping, location, and scale parameters of a Fréchet distribution (dotted line) to fit the measured data in (d). Reprinted from Dif-Pradalier et al. [68] (Copyright 2017 IOP) and Choi et al. [80] (Copyright 2024 IOP), respectively.
@@
The scaling of $\Delta$ with respect to $\rho_{*}=\rho_{s}/a$ has been of particular interest since it can be related to the global confinement scaling. Here, $\rho_{s}$ and $a$ are the ion sound gyroradius and the minor radius of plasma, respectively. In the GYSELA simulations [68], $\Delta$ does not seem to simply increase as $\rho_{*}^{-1}=a/\rho_{s}$ increases as shown in Figure 10(a). Instead, the most probable value has been fixed around $40~\rho_{s}$ for investigated values of $\rho_{*}$. This feature expects the favorable gyro-Bohm scaling. However, in Tore Supra experiments, $\Delta$ seems to increase with $\rho_{*}^{-1}$ as shown in Figure 10(b), implying that even with the avalanche regulation by the $E\times B$ staircase, the global confinement scaling can be like the Bohm scaling [77]. Further research would be required to address this inconsistency.
@@
The distribution of $\Delta$ was found to follow a long tail distribution called the Fréchet distribution in the GYSELA simulations, i.e. $P(\Delta)\sim P_{\text{Fréchet}}(\Delta)$ [68], and also in KSTAR experiments [80] as shown in Figures 10(c) and (d). The Fréchet distribution is one of the extreme value distributions, a distribution of maxima of samples when the samples follow a long tail distribution such as the power-law distribution. This Fréchet distribution can imply dual roles of avalanches, co-existing ballistic transport events (mixing process) as explained in reference [80]. They can destruct the $\mathbf{E}\times\mathbf{B}$ staircase but may also contribute to its formation and $\Delta$ determination [^3] The distribution of temperature perturbation size of avalanches ($S\sim\delta T^{2}$ [80]) was found to follow a power-law distribution ($P(S)\sim P_{\mathrm{power-law}}(S)$) [42]. Assume that (1) only a temperature perturbation beyond a large size threshold can trigger the formation of the staircase and (2) $\Delta$ depends on the size of the triggering temperature perturbation which is rare and whose size would be the maximum size of perturbations for some period. Then, the Fréchet distribution of $\Delta$ follows from its definition ($P(\Delta\sim\max\{S\})\sim P_{\text{Fréchet}}(\Delta)$). Some models have shown these behaviors, e.g., the threshold [86] and the turbulence spreading strength $\sim\Delta$ [87]. Also, note that there are minor temperature corrugations that form away from large avalanches without apparent interactions, and their $\Delta$ distribution does not seem to follow a Fréchet distribution [80]..
@@ 4 Summary and discussion
    There exist strong interactions between blobs/voids and $\mathbf{E}\times\mathbf{B}$ sheared flow. The turbulence spreading induced by blobs decreases as the flow shearing rate increases, while the turbulence spreading induced by voids is insensitive to the shearing rate. On the other hand, voids can excite drift waves and so drive intermittent zonal flow.
@@
The $\mathbf{E}\times\mathbf{B}$ staircase, or more generally mesoscopic transport barriers, represents a realization of inhomogeneous mixing in fusion plasmas. Various methods have been utilized to identify these elusive transport barriers in fusion core plasmas. As summarized in Table 3, they have a lifetime order of 0.1–1.0 ms with a long tail distribution (KSTAR), and a tread width of 10–100 $\rho_{s}$ (Tore Supra, HL-2A, TJ-II, W7-X) with a Fréchet distribution (KSTAR). Their strength seems to decrease with $\rho_{*}^{-1}$ (Tore Supra) or $\nu_{\mathrm{i}}$ (KSTAR), and they seem to be associated with the low-order rational surfaces (HL-2A, KSTAR, TJ-II, W7-X).
It may be worth discussing potentially complex effects of the low-order rational surfaces on plasma inhomogeneity or the $\mathbf{E}\times\mathbf{B}$ staircase. On the one hand, the low-order rational surface can be a region of enhanced mixing [88, 89]. The transport flux by more densely packed $m/n$ modes at the low-order rational surface can be relatively higher than that in adjacent regions, leading to the local flattening of profiles [88]. The low-order rational surface can be an initiation location of avalanche-like transport events [42, 89]. On the other hand, the generation of zonal flows out of turbulence self-interactions can be also more facilitated at the low-order rational surface [90] since eddies can bite their own tail more easily with the shorter parallel length [91]. The existence of rational surfaces can signify the ubiquitous inhomogeneous mixing (corrugated profiles) in magnetically confined toroidal plasmas.
According to the recent simulation [91] and theory [92, 93], the inhomogeneous self-generation of zonal flow at the rational surface can be more effective with the higher fraction of the fast ions, lowering the zonal flow generation threshold [92, 93]. This implies the stronger plasma inhomogeneity in future fusion experiments with significant energetic ions. Interestingly, avalanches and mesoscopic transport barriers in the electron channel were readily observed in the low-level-MHD phase of the KSTAR fast-ion regulated enhancement (FIRE) mode [94], characterized by the high fast ion fraction [95] The barriers seem to have a longer lifetime in the FIRE mode compared to the similar L-mode. If long stationary mesoscopic transport barriers could be realized in future experiments, it would be an attractive confinement regime which is free from disruptive MHD instabilities.
Although some progress has been made over the past decade in the experimental side, a comprehensive understanding of the physics of the $\mathbf{E}\times\mathbf{B}$ staircase remains elusive. A brief introduction of existing specific models on the $\mathbf{E}\times\mathbf{B}$ staircase is provided in Appendix.
In the physics literature, there are three classes of models which address staircase formation. These are (i) flux bistability, (ii) homogenization, and (iii) phase separation (spinodal decomposition). These are explained below.
@@
    Flux bistability depends on the coexistence of two different values of the local gradient at the same value of the total flux. Of course, the total flux is set by the sum of turbulent and collisional (neoclassical) fluxes, and the ratio of these two varies considerably. Bistability is described by the familiar S-curve for the highly nonlinear flux-gradient relation (see Figure 11(a)). This is familiar from studies of single barriers. However, a staircase is an array or lattice of (small) barriers, with regions of strong mixing interspersed between (see Figure 11(a)) [66, 96]. The classic example of a bistable mixing model of staircases is that of Balmforth, Llewellyn-Smith and Young (hereafter BLY) [96], which has yielded a bounty of insights into staircase structure and evolution. The BLY model aimed to explain the observations of density staircases in stable-but-buoyant mixing. The key feature of BLY is its use of two mixing lengths in an otherwise standard K-model approach. In BLY, the mixing scale $l$ is related to an imposed stirring scale $l_{ext}$ and an emergent scale $l_{oz}$, the Ozmidov scale [97], by $1/l=1/l_{ext}+1/l_{oz}$. The Ozmidov scale is dependent upon turbulence intensity and the density gradient, and is defined by the cross-over of the buoyancy frequency and the eddy turn over rate. The flux-gradient relation for the system exhibits a structure similar to the S-curve, discussed above. The density profile consists of an array of strong mixing zones interleaved by steep gradient layers (barriers).
@@
    Homogenization refers to the well-known tendency of eddies with differentially rotating flow to homogenize the profile of vorticity, scalar concentration, etc. within the eddy. Then, an array of tangential eddies, as shown in Figure 11(c), in a fluid with molecular diffusivity $D_{0}$ and large Peclet number $Pe=vl/D_{0}\gg 1$, naturally leads to inhomogeneous mixing [98, 99]. This is because mixing within the eddy is much faster than the slow diffusion through the separatrices between eddies. Thus, scalar concentration profiles naturally take the form of staircases, i.e. flat within the eddies and steep in the regions between them, as in Figure 11(d).
@@
    Phase separation [100] refers to the property of some mixtures which allows them to separate into domains of different constituent components (see Figure 11(e) [101]). Such a process also goes by the name of spinodal decomposition. Phase separation can lead to layering, as in Figure 11(f) [102], where a target pattern results, formed by alternating bonds of different phases.
Figure 11: (a) S-curve relating heat flux $Q$ and temperature gradient $\nabla T$. (b) Staircase profile of temperature $T$. (c) An array of vortex cells and (d) the averaged (in $y^{\prime}$ direction) scalar concentration profile $h(x^{\prime})$ in the form of the staircase. Reprinted from Ramirez and Diamond [99]. Copyright 2024 APS. (e) An illustration of phase separation in the binary fluid. Colors indicate the normalized component density contrast $\psi=(\rho_{A}-\rho_{B})/(\rho_{A}+\rho_{B})$. Reprinted from Fan et al. [101]. Copyright 2018 AIP. (f) Evolution of $\psi$ patterns from initial spirals to concentric annuli (staircase). Reprinted from Fan et al. [102]. Copyright 2017 APS.
The results of recent blobs/voids research can provide helpful insights into developing a better physical model for the $\mathbf{E}\times\mathbf{B}$ staircase (avalanches and zonal flow layers). Incorporating turbulence intensity flux and its dynamics into physical models of the $\mathbf{E}\times\mathbf{B}$ staircase is necessary. In fact, such models have been developed but have not yet been exercised effectively. In particular, the length scale for the turbulence intensity can differ from that for the transport flux, adding a scale to the bistable mixing model [87]. This is plausible, since research on coherent structures [7] indicates a departure of the spreading flux from the particle flux. The scale of the spreading flux effectively sets the scale of the turbulence envelope, which in turn determines the zonal flow.
Figure 4 shows that the symmetry breaking of turbulence intensity flux, carried by the coherent structures such as blobs/voids, yields a net turbulence spreading and forms a region of enhanced mixing. On the other hand, recent studies [6, 65] introduced in item 2.4 of Section 2 show that the blobs/voids can contribute to the zonal flow generation in core plasmas as they propagate across the plasma boundary[103, 104, 9]. The resulting sheared flow layer forms a region of suppressed mixing. The correlated growth of zonal flows at distant regions might be mediated by turbulence spreading of momentum flux between zonal flow layers [105, 106]. Therefore, relating the coherent structures and their propagation to the formation or sustainment of the $\mathbf{E}\times\mathbf{B}$ staircase should not be surprising. Indeed, some relevant observations were already reported, i.e. the triggering of temperature staircase by the excessive localized heating [80].
A key question for the $\mathbf{E}\times\mathbf{B}$ staircase is the scale selection problem – i.e. what sets the separation between zonal flow layers? There are several possibilities:
@@
    Bistable mixing (two scales model). Here, the emergent scale adjusts to define the layer separation. This is an outcome of the nonlinear flux-gradient relation. In the case of drift-Rossby turbulence, the Rhines scale [107] is selected (see Appendix).
@@
Further research – experimental, theoretical and computational – is necessary to sort out which mechanism is relevant. Experimental research includes, but is not limited to, (1) the multi-machine analysis, (2) the development of a stable staircase scenario (possibly with the help of energetic particles [92, 93] or light impurities [108]), and (3) the development of fine-scale flow or potential diagnostics, which enables the detailed analysis of inhomogeneous mixing process (see Section 2), in the core plasma.
@@
4.1 Production ratio for gyrokinetic ITG model
The basic equation is the gyrokinetic equation and the relevant intensity is $\langle g^{2}\rangle$, where $g$ is the non-adiabatic distribution function. We briefly sketch a derivation for the simplified case of electrostatic ion drift kinetics in a straight field. The basic equation is:
$\frac{\partial g}{\partial t}+\mathbf{v}_{\perp}\cdot\nabla g=\frac{|e|}{T_{\mathrm{i}}}\frac{\partial\phi}{\partial t}\langle f\rangle+v_{*\mathrm{i}}(E)\frac{\partial}{\partial y}\left(\frac{|e|\phi}{T_{\mathrm{i}}}\right)\langle f\rangle$ (1)
Here, $\langle f\rangle$ is the inhomogeneous mean distribution and $v_{*\mathrm{i}}(E)$ is the energy-dependent full diamagnetic velocity (including temperature gradient dependent piece).
Multiplying by g gives:
$\frac{\partial\langle g^{2}\rangle}{\partial t}+\nabla\cdot\langle\mathbf{v}_{\perp}gg\rangle=\frac{|e|}{T_{\mathrm{i}}}\langle f\rangle\left[\left\langle\frac{\partial\phi}{\partial t}g\right\rangle+v_{*\mathrm{i}}(E)\left\langle\frac{\partial\phi}{\partial y}g\right\rangle\right]$ (2)
The kinematic production ratio then is the comparison of the velocity integrated second term to that for the third term.
Integrating over velocity space gives (for a steady state):
$\nabla\cdot\int d^{3}v\langle\mathbf{v}_{\perp}gg\rangle=\int d^{3}v\frac{|e|}{T_{\mathrm{i}}}\langle f\rangle\left[\left\langle\frac{\partial\phi}{\partial t}g\right\rangle+v_{*\mathrm{i}}(E)\left\langle\frac{\partial\phi}{\partial y}g\right\rangle\right]$ (3)
The production ratio can be given by
$PR=d\left(\int d^{3}v\langle\tilde{v}_{\perp}gg\rangle\right)/\mathcal{R}$ (4)
where $\mathcal{R}=\int_{r-\delta}^{r+\delta}dr\int d^{3}v\frac{|e|}{T_{\mathrm{i}}}\langle f\rangle\left[\left\langle\frac{\partial\phi}{\partial t}g\right\rangle+v_{*\mathrm{i}}(E)\left\langle\frac{\partial\phi}{\partial y}g\right\rangle\right]$ and $d(F)\equiv F(r+\delta)-F(r-\delta)$. Note that $PR$ is still the ratio of the differential of the spreading flux in phase space to the integrated production due to phase space relaxation. In practice, the kinetic $PR$ here can be calculated only via simulation.
4.2 Models on the $\mathbf{E}\times\mathbf{B}$ staircase